C2H2 can be produced by the reaction of calcium carbide with water according to the following equation

CaC2(s)+ H2O(I) = C2H2 (g)+Ca(OH)2(aq) if 48.6g of calcium carbide is reacted at STP, what is the volume of the acetylene produced?


1
Expert's answer
2021-02-09T03:39:45-0500

Solution:

The balanced chemical equation:

CaC2(s) + 2H2O(I) = C2H2(g) + Ca(OH)2(aq)

According to the equation above: n(CaC2) = n(C2H2)


Moles of CaC2 = Mass of CaC2 / Molar mass of CaC2

The molar mass of CaC2 is 64.099 g mol-1.

Hence,

n(CaC2) = 48.6 g / 64.099 g mol-1 = 0.7582 mol


n(C2H2) = n(CaC2) = 0.7582 mol


At STP, one mole of any gas occupies a volume of 22.4 L.

Thus 0.7582 mol of C2H2 occupies:

(0.7582 mol × 22.4 L) / 1 mol = 16.98 L of C2H2

V(C2H2) = 16.98 L = 17.0 L


Answer: 17.0 L of the acetylene (C2H2) produced.